# MoE-FFD: Mixture of Experts for Generalized and Parameter-Efficient Face Forgery Detection

### v2

@@ 4.1 Implementation Details
We apply the popular MTCNN face detector [68] to crop the face regions. The proposed framework is implemented on the Pytorch [69] platform. The model is trained using Adam optimizer [70] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate of gating network and other trainable parameters as 1e-4 and 3e-5, respectively. The loss weigh $\lambda$ in Eq. (5) is set as 1. We train the model for 20 epochs on one single 3090 GPU with batch size 32. We follow the official dataset split strategy in [6] for fair comparison.
TABLE II: Cross-manipulation detection AUC on the unseen manipulation technique (FF++ C23 dataset).
Method #Params DF FF FS NT AVG
@@ 4.2 Datasets and Evaluation Metrics
The experiments encompass the following six datasets for training and testing: FaceForensics++ (FF++) [72], CelebDF-v2 (CDF) [27], WildDeepfake (WDF) [73], DeepFake Detection Challenge Preview (DFDC-P) [74], DeepFakeDetection (DFD) [75], and DeepForensics-1.0 (DFR) [76]. FF++ is a widely used dataset in face forgery detection, which includes four face manipulation types: Deepfakes (DF) [77], Face2Face (FF) [78], FaceSwap (FS) [79], and NeuralTextures (NT) [80]. The remaining datasets represent recent five Deepfake datasets with diverse environment variables and various video qualities. We perform cross-dataset and cross-manipulation evaluations to examine the models’ generalization capability. Additionally, we measure the model’s robustness by applying it to perturbed face images with various perturbation types and severity levels. Consistent with prior arts, we adopt Area under the ROC Curve (AUC) and Equal Error Rate (EER) as evaluation metrics. AUC measures the Area under the Receiver Operating Characteristic (ROC) curve, while EER denotes the False Positive Rate (FPR) that equals to the True Positive Rate (TPR).
TABLE III: Cross-manipulation detection AUC on the remaining three unseen manipulation techniques (FF++ C23 dataset).
Methods Train DF FF FS NT AVG
@@
Ours NT 80.02 73.02 51.94 98.70 68.33
@@ 4.3 Comparison with State-of-the-Arts
In this subsection, we perform cross-dataset evaluations to evaluate the models’ generalization capability. In practical scenarios, the trained models are vulnerable to unseen domains, leading to significant performance degradations. We train all face forgery detectors on the FF++ (C23) dataset and directly evaluate them on five unseen Deepfake datasets: CDF, WDF, DFDC-P, DFD, and DFR. Table I presents the frame-level and video-level detection results, divided into top and bottom sections. Video-level detection results highlighted in blue color are obtained by averaging all frame scores within each video. We bold the best results and underline the second-best results. Notably, MoE-FFD achieves the highest detection performance on four datasets in both frame-level and video-level detection. Additionally, MoE-FFD’s performance is on par with CFM on the DFD dataset. Moreover, MoE-FFD outperforms CFM by a clear margin in average performance. The average frame-level AUC improves by 2.86%, rising from 83.94% to 86.78%. Similarly, MoE-FFD also outperforms others in video-level detection on average. Our MoE-FFD’s generalizability improvements over existing methods offers valuable insights into the effectiveness of the proposed approach.
Figure 4: Robustness to various common perturbations at five severity levels: Gaussian blur, pink noise, white noise, and block wise.
@@
Existing face forgery detectors often struggle to handle emerging manipulation techniques. With the rapid development of AIGC, more sophisticated manipulation techniques continue to emerge, posing challenges for existing detection models. Ensuring model generalizability to unseen forgeries is crucial for real-world applications. In this study, we conduct cross-manipulation experiments involving four forgery techniques: Deepfakes (DF), Face2Face (FF), FaceSwap (FS), and NeuralTextures (NT). Table II presents the results, where models trained on three manipulation types are tested on the remaining one. MoE-FFD exhibits the best average detection results. Compared to the ViT-B baseline, the proposed method achieves an average AUC enhancement of 18.5%, going from 62.3% to 80.8%. In Table III, we further examine models trained on one manipulation type and tested across the other three. AVG denotes the average AUC score across three cross-manipulation trials. MoE-FFD outperforms the baseline in both intra- and cross-manipulation evaluations. Remarkably, MoE-FFD achieves around 99% AUC scores in each intra-manipulation evaluation, highlighting its outstanding detection accuracy in intra-domain evaluations. When compared to the state-of-the-art RECCE method, our MoE-FFD consistently achieves superior average cross-manipulation results across all four evaluation settings. These findings suggest that MoE-FFD is generalized to unseen manipulations.
TABLE IV: Ablation experiment on the designed MoE modules.
@@
Images and videos transmitted online always undergo various perturbations that erase forgery cues within the image/video contents [81]. As such, the detection performance of existing models significantly drops on distorted data. In this work, we introduce common perturbations to measure the model’s robustness, including Gaussian blurring, pink noise, white noise, and blockwise. Each perturbation type further involves five severity levels to mimic diverse real-world conditions. Note that we do not apply any data augmentations during training, such that the tested perturbations are totally unseen for our model. As shown in Fig. 4, the detectors’ AUC detection performance consistently deteriorates with the increasing severity level. However, our proposed MoE-FFD exhibits significantly greater resilience to most perturbations compared to previous methods. This indicates that the proposed MoE-FFD effectively captures the inherent forgery clues. Notably, MoE-FFD achieves substantial improvements in robustness compared to the ViT-B baseline, demonstrating that the designed MoE learning scheme and the PEFT modules greatly enhance the model’s robustness against image perturbations.
@@
TABLE V: Cross-dataset detection performance on different ViT backbones.
@@ 4.4 Ablation Experiments
The proposed MoE-FFD integrates one LoRA layer and one Adapter layer with each ViT block. The LoRA layer is designed to learn forgery-specific parameters for subsequent attention mechanisms, capturing long-range interactions within the input faces. Meanwhile, the Convpass Adapter introduces forgery local priors into the plain ViT model. To study the effectiveness of the designed PEFT modules, we report the cross-dataset detection results in Table IV. Utilizing MoE LoRA and MoE Adapter significantly improves the model generalizability across all datasets compared to the vanilla ViT-B backbone. This enhancement can be attributed to the LoRA and Adapter layers effectively retaining the ImageNet knowledge while adaptively learning forgery features. Furthermore, the combination of LoRA and Adapter layers allows the model to capture both long-range interactions and local forgery cues, further enhancing its face forgery detection performance.
@@
Flexibility is crucial in real-world applications to address the complexities of practical scenarios. Deployment devices with varying computational resources may necessitate different detection models. Therefore, we integrate the proposed MoE-FFD into ViT-Tiny, ViT-Small, and ViT-Large models to assess the flexibility of our approach. MoE-FFD can be readily inserted to other vision transformer backbones in a plug-and-play manner. The cross-dataset results are presented in Table V. Compared with directly finetuning the vanilla ViT backbones, MoE-FFD significantly reduces the training parameters. As the model size increases, the model generalizability consistently enhances. Additionally, significant performance improvements are observed across different ViT backbones by using MoE-FFD. Specifically, when we assemble MoE-FFD with the ViT-Large backbone, the model achieves 13.08%, 3.05%, and 7.04% AUC boosts on CDF, DFDC-P, and WDF datasets, with only $\sim$13.6trainable parameters. This, in turn, demonstrates the flexibility and effectiveness of our method.
TABLE VI: Impacts of the matrix rank of LoRA layers and the effectiveness of MoE.
@@
In this work, the proposed MoE is designed to dynamically select the optimal Top-1 expert for face forgery detection. In Table VI, we compare the cross-dataset detection performance of the proposed MoE-FFD with individual LoRA experts of varying ranks. Notably, different datasets exhibit varying optimal LoRA ranks. The model reaches the best AUC scores on CDF, DFDC-P, and WDF datasets with the LoRA rank of 64, 96, and 48, respectively. MoE dynamically selects the LoRA rank, facilitating the model to search the optimal feature space for each input query. In Table VI, MoE consistently outperforms individual LoRA experts on all three datasets. Similarly, we investigate the impact of different Adapters in Table VII. As introduced in Sec. 3.3.2, the designed adapters tend to expose different local artifacts of input faces. MoE smartly searches the best local feature extractor for each input real/fake face, thereby achieving the superior results compared to using a single Adapter. It should be noted that the MoE approach only introduces negligible additional activated parameters, demonstrating that the performance improvements stem from the proposed MoE learning scheme, instead of the model scaling up.
@@
TABLE VII: Impacts of different designed Adapters and the effectiveness of MoE.
@@
TABLE VIII: MoE v.s. Multi-Experts.

### v3

@@ 4.1 Implementation Details
We apply the popular MTCNN face detector [83] to crop the face regions. The proposed framework is implemented on the Pytorch [84] platform. The model is trained using Adam optimizer [85] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate of gating network and other trainable parameters as 1e-4 and 3e-5, respectively. The loss weight $\lambda$ in Eq. (5) is set as 1. We train the model for 20 epochs on one single 3090 GPU with batch size 32. We follow the official dataset split strategy in [6] for fair comparison. In the proposed MoE-FFD framework, the ViT-Base model pre-trained on ImageNet-21K [86] serves as the backbone. It consists of 12 transformer blocks, each comprising a 12-head self-attention layer and a multi-layer perceptron (MLP) layer.
TABLE II: Cross-dataset evaluation on DF40 (CDF).
Method FS (CDF) FR (CDF) EFS (CDF) AVG
AUC EER AUC EER AUC EER AUC EER
SBI[33] 70.01 35.20 48.90 50.98 65.61 39.21 61.50 41.79
CFM[6] 78.65 28.73 69.11 35.96 71.49 34.33 73.08 33.00
DCL[34] 72.67 35.04 62.88 41.01 63.18 41.20 66.24 39.08
ViT-B [19] 67.78 37.19 55.65 46.57 60.50 42.64 61.31 42.13
MoE-FFD 82.11 26.97 66.54 37.92 78.17 29.78 75.60 31.55
TABLE III: Cross-manipulation detection AUC on the unseen manipulation technique (FF++ C23 dataset).
Method #Params DF $F$ FS NT AVG
@@ 4.2 Datasets and Evaluation Metrics
The experiments encompass the following six datasets for training and testing: FaceForensics++ (FF++) [88], CelebDF-v2 (CDF) [27], WildDeepfake (WDF) [89], DeepFake Detection Challenge Preview (DFDC-P) [90], DeepFakeDetection (DFD) [91], and DeepForensics-1.0 (DFR) [92]. FF++ is a widely used dataset in face forgery detection, which includes four face manipulation types: Deepfakes (DF) [93], Face2Face (FF) [94], FaceSwap (FS) [95], and NeuralTextures (NT) [96]. The remaining datasets represent recent five Deepfake datasets with diverse environment variables and various video qualities. We perform cross-dataset and cross-manipulation evaluations to examine the models’ generalization capability. Specifically, CDF is a high-quality DeepFake dataset specifically focused on celebrity identities. WDF is a web-crawled DeepFake dataset that contains videos with diverse scenes and forgery methods. DFDC-P is a challenging dataset that contains deepfake videos generated from various scenarios using two face manipulation methods. DFD is a comprehensive DeepFake dataset comprising over 3,000 manipulated videos involving 28 actors across various scenes. DFR is created by using real videos from FF++ with an innovative end-to-end face swapping framework. In addition, we examine our method’s effectiveness on the recently released DF40 benchmark [97] with 40 forgery types. It comprises 10 face swapping (FS) methods, 13 face reenactment (FR) methods, 12 entire face synthesis (EFS) methods, and 5 face editing (FE) methods. Real data is sourced from the FF++ [88] and CDF [27] datasets, and the corresponding fake data is generated to construct domain-specific subsets for each manipulation type, such as FS (FF++) and FS (CDF). Furthermore, we measure the model’s robustness by applying it to perturbed face images with various perturbation types and severity levels. Consistent with prior arts, we adopt Area under the ROC Curve (AUC) and Equal Error Rate (EER) as evaluation metrics. AUC measures the Area under the Receiver Operating Characteristic (ROC) curve, while EER denotes the False Positive Rate (FPR) that equals to the True Positive Rate (TPR).
TABLE IV: Cross-manipulation detection AUC on the remaining three unseen manipulation techniques (FF++ C23 dataset).
Methods Train DF $F$ FS NT AVG
@@
MoE-FFD NT 80.02 73.02 51.94 98.70 68.33
TABLE V: Cross-manipulation evaluation on DF40 (FF++). (AUC)
Training Set Model Testing Set
FS (FF++) FR (FF++) EFS (FF++) AVG (FF++)
FS (FF++) Xception [9] 0.991 0.892 0.810 0.898
CLIP [98] 0.996 0.908 0.837 0.914
SRM [11] 0.988 0.867 0.703 0.853
SPSL [99] 0.987 0.849 0.735 0.857
RECCE [35] 0.991 0.855 0.758 0.868
RFM [100] 0.992 0.884 0.821 0.899
ViT-B [19] 0.983 0.877 0.771 0.877
MoE-FFD 0.990 0.979 0.904 0.957
FR (FF++) Xception [9] 0.838 0.996 0.670 0.835
CLIP [98] 0.932 0.999 0.798 0.910
SRM [11] 0.893 0.998 0.698 0.863
SPSL [99] 0.901 0.998 0.695 0.865
RECCE [35] 0.865 0.997 0.716 0.859
RFM [100] 0.892 0.999 0.776 0.889
ViT-B [19] 0.764 0.986 0.645 0.798
MoE-FFD 0.933 0.999 0.866 0.932
EFS (FF++) Xception [9] 0.665 0.807 0.999 0.824
CLIP [98] 0.688 0.889 0.999 0.859
SRM [11] 0.596 0.776 0.999 0.790
SPSL [99] 0.659 0.811 0.999 0.823
RECCE [35] 0.691 0.801 0.999 0.830
RFM [100] 0.653 0.795 0.999 0.816
ViT-B [19] 0.550 0.718 0.998 0.755
MoE-FFD 0.737 0.854 0.999 0.863
@@ 4.3 Comparison with State-of-the-Arts
4.3.1 Baseline Models.
This paper incorporates 14 representative baseline detectors from top journals and conferences, including three data-driven architectures and 11 state-of-the-art image forgery detectors:
Face X-ray [4] detects face forgeries by revealing blending boundaries indicative of image composition from different sources.
GFF [11] leverages high-frequency noise information for face forgery detection, which is less affected by manipulation-specific textures and more indicative of tampered regions.
LTW [73] uses meta-weight learning to assign adaptive weights to training samples from different domains, mitigating domain-specific biases and enhancing model generalizability.
F2Trans-S [26] is a high-frequency fine-grained transformer network for face forgery detection that enhances generalization and robustness by leveraging fine-grained manipulation traces in both spatial and frequency domains.
SBI [33] encourages the model to learn more general and manipulation-agnostic features by blending pseudo source and target images from single pristine images.
DCL [34] enhances generalization by constructing positive and negative pairs for contrastive learning at both inter- and intra-instance levels.
RECCE [35] focuses on learning compact and generalizable representations of genuine faces rather than specific forgery patterns. By combining multi-scale reconstruction learning with classification learning, RECCE effectively enhances the model robustness and generalization.
SFDG [74] enhances model’s generalization capability by modeling relation-aware features across spatial and frequency domains.
UCF [75] is a disentanglement-based framework for deepfake detection that improves generalization by isolating and leveraging common forgery features while minimizing reliance on forgery-irrelevant or method-specific patterns.
LSDA [76] mitigates overfitting to method-specific artifacts and enables a more robust decision boundary by simulating intra- and inter-forgery variations, thereby achieving superior performance across diverse datasets.
CFM [6] enhances face forgery detection by mining critical forgery features through prior knowledge-agnostic data augmentation, fine-grained relation learning, and a progressive learning controller.
@@
In this subsection, we conduct cross-dataset experiments to evaluate the models’ generalization capability. In practical scenarios, the trained models are vulnerable to unseen domains, leading to significant performance degradations. We train all face forgery detectors on the FF++ (C23) dataset and directly test them on five unseen Deepfake datasets: CDF, WDF, DFDC-P, DFD, and DFR. Table I presents the frame-level and video-level detection results, divided into top and bottom sections. Video-level detection results presented in the bottom table are obtained by averaging all frame scores within each video. We bold the best results and underline the second-best results. Notably, MoE-FFD achieves the highest detection performance on four datasets in both frame-level and video-level detection. Additionally, MoE-FFD’s performance is on par with CFM on the DFD dataset. Moreover, MoE-FFD outperforms CFM by a clear margin in average performance. The average frame-level AUC improves by 2.86%, rising from 83.94% to 86.78%. Similarly, MoE-FFD also outperforms others in video-level detection on average. Our MoE-FFD’s generalizability improvements over existing methods offers valuable insights into the effectiveness of the proposed approach.
We present the cross-dataset evaluation results on the DF40 dataset in Table II. The listed models are trained on FF++ (C23) and evaluated on FS (CDF), FR (CDF), and EFS (CDF), which are generated from the CDF dataset. It can be observed that MoE-FFD exhibits superior generalization performance on the unforeseen datasets, further validating its effectiveness for face forgery detection.
Unlike previous arts, MoE-FFD adopts a parameter-efficient fine-tuning strategy that preserves pretrained knowledge while dynamically adapting to forgery-specific features. Furthermore, MoE-FFD effectively scales up the network’s capacity by introducing multiple specialized forgery experts, enabling the model to dynamically select the most appropriate expert when processing unseen facial forgeries.
Figure 4: Robustness to various common perturbations at five severity levels: Gaussian blur, pink noise, white noise, block wise, JPEG compression, and Gaussian noise.
@@
Existing face forgery detectors often struggle to handle emerging manipulation techniques. With the rapid development of AIGC, more sophisticated manipulation techniques continue to emerge, posing challenges for existing detection models. Ensuring model generalizability to unseen forgeries is crucial for real-world applications. In this study, we conduct cross-manipulation experiments involving four forgery techniques: Deepfakes (DF), Face2Face (FF), FaceSwap (FS), and NeuralTextures (NT). Table III presents the results, where models trained on three manipulation types are tested on the remaining one. MoE-FFD exhibits the best average detection results. Compared to the ViT-B baseline, the proposed method achieves an average AUC enhancement of 18.5%, going from 62.3% to 80.8%. In Table IV, we further examine models trained on one manipulation type and tested across the other three. AVG denotes the average AUC score across three cross-manipulation trials. MoE-FFD outperforms the baseline in both intra- and cross-manipulation evaluations. Remarkably, MoE-FFD achieves around 99% AUC scores in each intra-manipulation evaluation, highlighting its outstanding detection accuracy in intra-domain evaluations. When compared to the state-of-the-art RECCE method, our MoE-FFD consistently achieves superior average cross-manipulation results across all four evaluation settings. In addition, we further perform cross-manipulation evaluations on the DF40 benchmark. Following the official protocol, the model is trained on a single manipulation type (i.e., FS, FR, or EFS) and evaluated across all manipulation types within the FF++ domain. As shown in Table V, the proposed MoE-FFD consistently achieves the highest average AUC scores across various cross-manipulation settings, demonstrating superior performance compared to other SOTA methods. These findings suggest that MoE-FFD is generalized to unseen manipulations.
TABLE VI: Ablation experiment on the designed MoE modules.
@@
Images and videos transmitted online always undergo various perturbations that erase forgery cues within the image/video contents [101]. Existing Deepfake detectors always suffer from performance drops on distorted data. In this work, we introduce common perturbations to evaluate the model’s robustness, including Gaussian blurring, pink noise, white noise, block-wise, JPEG compression, and Gaussian noise. Each perturbation type includes five severity levels to mimic diverse real-world conditions. Note that we do not apply any data augmentations during training, such that the tested perturbations are totally unseen for our model. As shown in Fig. 4, the detectors’ AUC detection performance consistently deteriorates with the increasing severity level. Nevertheless, our proposed MoE-FFD demonstrates higher resilience to most perturbations compared to existing methods. Notably, it achieves substantial robustness gains over the ViT-B baseline, demonstrating the effectiveness of the proposed MoE learning strategy and PEFT modules in improving robustness against image perturbations. This further suggests that MoE-FFD can effectively capture intrinsic forgery features.
@@
TABLE VII: Cross-dataset detection performance on different ViT backbones.
@@ 4.4 Ablation Experiments
The proposed MoE-FFD integrates one LoRA layer and one Adapter layer with each ViT block. The LoRA layer is designed to learn forgery-specific parameters for subsequent attention mechanisms, capturing long-range interactions within the input faces. Meanwhile, the Convpass Adapter introduces forgery local priors into the plain ViT model. To study the effectiveness of the designed PEFT modules, we report the cross-dataset detection results in Table VI. Utilizing MoE LoRA and MoE Adapter significantly improves the model generalizability across all datasets compared to the vanilla ViT-B backbone. This enhancement can be attributed to the LoRA and Adapter layers effectively retaining the ImageNet knowledge while adaptively learning forgery features. Furthermore, the combination of LoRA and Adapter layers allows the model to capture both long-range interactions and local forgery cues, further enhancing its face forgery detection performance.
@@
Flexibility is crucial in real-world applications to address the complexities of practical scenarios. Deployment devices with varying computational resources may necessitate different detection models. Therefore, we integrate the proposed MoE-FFD into ViT-Tiny, ViT-Small, and ViT-Large models to assess the flexibility of our approach. MoE-FFD can be readily inserted to other vision transformer backbones in a plug-and-play manner. The cross-dataset results are presented in Table VII. Compared with directly finetuning the vanilla ViT backbones, MoE-FFD significantly reduces the training parameters. As the model size increases, the model generalizability consistently enhances. Additionally, significant performance improvements are observed across different ViT backbones by using MoE-FFD. Specifically, when we assemble MoE-FFD with the ViT-Large backbone, the model achieves 13.08%, 3.05%, and 7.04% AUC boosts on CDF, DFDC-P, and WDF datasets, with only $\sim$13.6% trainable parameters. This, in turn, demonstrates the flexibility and effectiveness of our method.
TABLE VIII: Impacts of the matrix rank of LoRA layers and the effectiveness of MoE.
@@
In this work, the proposed MoE is designed to dynamically select the optimal Top-1 expert for face forgery detection. In Table VIII, we compare the cross-dataset detection performance of the proposed MoE-FFD with individual LoRA experts of varying ranks. Notably, different datasets exhibit varying optimal LoRA ranks. The model reaches the best AUC scores on CDF, DFDC-P, and WDF datasets with the LoRA rank of 64, 96, and 48, respectively. MoE dynamically selects the LoRA rank, facilitating the model to search the optimal feature space for each input query. In Table VIII, MoE consistently outperforms individual LoRA experts on all three datasets. Similarly, we investigate the impact of different Adapters in Table IX. As introduced in Sec. 3.3.2, the designed adapters tend to expose different local artifacts of input faces. MoE smartly searches the best local feature extractor for each input real/fake face, thereby achieving the superior results compared to using a single Adapter. It should be noted that the MoE approach only introduces negligible additional activated parameters, demonstrating that the performance improvements stem from the proposed MoE learning scheme, instead of the model scaling up.
@@
TABLE IX: Impacts of different designed Adapters and the effectiveness of MoE.
@@
TABLE X: Ablation experiment on the selection of Top-k.
Top-k CDF DFDC-P WDF
AUC EER AUC EER AUC EER
k=3 84.01 24.24 80.00 28.27 78.84 29.51
k=2 85.29 23.62 79.49 28.48 79.44 28.82
k=1 86.69 22.06 80.83 26.67 80.64 27.11
4.4.4 Impacts of the Number of Activated Experts.
To further investigate the impact of the number of activated experts in the Top-k noisy gating mechanism, we conducted an ablation study, as presented in Table X. The results show that selecting the Top-1 expert consistently yields the highest AUC scores across all unseen datasets, demonstrating its superior generalization capability. In contrast, activating multiple experts may introduce noise from less confident or less relevant ones. Given that face forgery detection is a binary classification task, incorporating weaker experts with the accurate Top-1 expert can dilute informative signals [58], thereby deteriorating the detection performance.
4.4.5 MoE v.s. Multi-Experts.
[…]
4.4.6 Effectiveness of MoE Loss.
The gating network often exhibits a tendency to consistently assign large weights to only a few experts [58], resulting in overfitting problems. To address this issue, we introduce an MoE loss component aimed at encouraging equal importance among all experts. This regularization also prevents the model from getting trapped in local optima. In this subsection, we examine the impacts of the proposed $L_{moe}$. Table XII presents the cross-dataset evaluation performance with different values of the $L_{moe}$ loss weight $\lambda$ in Eq. (5), where $\lambda=0$ represents no MoE loss applied. We observe that the use of $L_{moe}$ effectively mitigates the gate overfitting problem and consistently boosts model generalizability. Furthermore, the model achieves the best generalizability with $\lambda=1$.
4.5 Visualization Results
TABLE XI: MoE v.s. Multi-Experts.